# Enabling Regional Explainability by Automatic and Model-agnostic Rule Extraction

### v2

@@ Abstract
Explainable Artificial Intelligence (XAI) [1] has garnered increasing attention in recent years, driven by the widespread adoption of deep learning models. XAI generally aims to enable professionals to understand why a model has made a specific prediction (decision) and/or what a model has learned from the data. Rule extraction (also known as rule learning) [2, 3] is a field in XAI that translates a model’s knowledge through a set of logical rules, such as IF-THEN statements, to obtain explainability of a model’s reasoning. Such rule-based explanations can offer interpretations that mimic human experts’ reasoning in solving knowledge-intensive problems [4]. This type of interpretations are particularly desirable in applications with tabular data, such as drug discovery [5] and healthcare [6].
@@
In practice, we often want to interpret a model’s analysis over a specified subgroup of the data. For instance, interpreting the patterns captured by a model for diagnosing a specific disease is highly desirable in clinical applications (Figure 1(a)). Data samples labelled with the disease must share similar patterns that do not exist outside the specific subgroup. However, the subgroups of interest usually have sporadic occurrences in many real-world scenarios, including most applications that prefer rule explanations, such as disease diagnosis, drug discovery, and financial fraud detection [7]. Existing methods in rule extraction focus on global explainability, often compromising regional explainability of minor subgroups of data to achieve better overall performance across the entire data distribution.
@@
(a) An example scenario of applying regional rule extraction for diabetes prediction. The data region of interest is the region that contains samples predicted as positive diabetes cases by the model. The results are from our experiments on a diabetes prediction task [8]. In the extracted rules, Hemoglobin A1c (HbA1c) is a measure of a person’s average blood sugar level over the past 2-3 months.
@@
(b) The workflow of our methods. Algorithm 1 is optional, depending on the dimensionality of the data.
Figure 1: An overview of our method, illustrating the regional rule extraction and depicting the process pipeline.
Regional rule extraction
@@
To better interpret data patterns for a specified subgroup, we propose a novel method for regional rule extraction, which aims to extract optimal rules from a specified data region. Formally, the outcome of regional rule extraction from a model is in the form of IF-THEN statements: IF $x\in\mathbb{X}$, THEN $y\in\mathbb{Y}$, here $\mathbb{X}$ denotes a subspace of a multivariate variable $x$ confined by a set of rules, e.g. value ranges of several features. $\mathbb{Y}$ denotes a subset of a target variable $y$. For instance, when extracting rules from a classifier, $y$ often refers to the model’s prediction and $\mathbb{Y}$ denotes a class of interest (e.g. positive class in binary classification).
@@
Compared to global rule extraction, regional rule extraction does NOT generate rules for the data region $y\notin\mathbb{Y}$. It is expected to gain two advantages through this reduction: i) obtaining more accurate and generalised rules for the specified data region $y\in\mathbb{Y}$, and ii) providing more precise control on key properties of the rules (e.g. number of rules, number of samples that satisfy these rules) for the specified data region.
A relevant topic in rule extraction is sequential covering, which iteratively learns a rule list for a data subgroup at each step to create a decision set that covers the entire dataset [9]. In sequential covering, the method applied to learn the rule list for a specified subgroup often involves a global rule extraction method, such as decision trees. The learned rules at each step then become the decision path to the purest node of the specified subgroup in a decision tree. To this end, regional rule extraction can be directly applied at each step of sequential covering to extract rules for specific data subgroups.
Implementation challenges
Conventional ways to extract rules for data subgroups or the entire dataset often utilise various search strategies, such as hill-climbing, beam search, and exhaustive search [9], as the objective is to search the best subspace of features that significantly elevates the occurrence of data of interest. Due to the vastness of the search space, these strategies often require predefined discretisation for numerical features,
For instance, association rule classifiers [10, 11] use association rule mining [12, 13] to transform rule extraction into finding frequent item sets – groups of items that frequently appear together in training samples. Consequently, numerical features must be transformed to discretised items before this process. Similarly, KDRuleEx [14] was developed to extract decision tables using categorical features. Moreover, Falling rule list [15], Scalable Bayesian Rule Lists (SBRL) [16], and Bayesian Rule Sets (BRS) [17] are probabilistic rule extraction methods that use distributions of discrete variables for modelling rule generation.
@@
However, predefined discretisation is often neither applicable nor desirable in real-world applications. Besides, when features are discretised individually, meaning the discretisation considers only the marginal distributions of features rather than their joint distributions, there is an implicit assumption that these features are independent – a condition not valid in many real-world scenarios. Moreover, the computational cost of automatic discretisation may increase exponentially when considering dependencies among a large number of features.
Decision Tree (DT) classifier [18, 19] is the most popular approach applied in rule extraction for neural networks due to its flexibility and accessibility [20, 21, 22]. The most attractive advantage of decision trees is their accessibility: they do not require predefined discretisation for numerical features. However, decision trees are not well-suited for regional rule extraction as: i) the selection of binary splits may still be affected by other data regions that are not of interest, especially in multi-mode scenarios; ii) the fact that they scan every observed value of a feature to find an optimal split, leading to a tendency to overfit the minimum support criterion on top features.
Automatic and Model-agnostic Regional Rule Extraction
@@
    we enhance regional rule extraction for black-box machine learning models, offering more accurate and generalised rules for underrepresented data regions than global rule extraction, and also allowing for local rule extraction for a given sample;
@@
    we enable automatic rule generation with numerical features, without requiring predefined discretisation or making any assumption about feature distributions;
@@
The workflow of our methods is illustrated in Figure 1(b). To the best of our knowledge, this is the first attempt to specifically address regional rule extraction with automated rule generation in the literature of XAI.
Results

### v3

@@ I Introduction
Explainable Artificial Intelligence (XAI) [1] has garnered increasing attention in recent years, driven by the widespread adoption of deep learning models. XAI generally aims to enable professionals to understand why a model has made a specific prediction (decision) and/or what a model has learned from the data. For these purposes, prior works in XAI often focus on producing either local or global XAI [2]. In global XAI, Rule extraction (also known as rule learning) [3, 4] is a field that translates a model’s knowledge through a set of logical rules, such as IF-THEN statements, to obtain explainability of a model’s reasoning. Such rule-based explanations can offer interpretations that mimic human experts’ reasoning in solving knowledge-intensive problems [5] and provide insights into key feature interactions in high dimensional spaces [6]. This type of interpretations are particularly desirable in knowledge-based applications of machine learning, such as drug discovery [7] and healthcare [8].
@@
In practice, we often want to interpret a model’s knowledge for identifying a specified subgroup of the data. For instance, interpreting the patterns captured by a model for diagnosing a specific disease is highly desirable in clinical applications (Figure 1). Data samples labeled with the disease must share similar patterns that do not exist outside the specific subgroup. However, the subgroups of interest usually have sporadic occurrences in many real-world scenarios, as a minor class in classifications tasks. Most applications that prefer rule explanations, such as disease diagnosis, drug discovery, and financial fraud detection [9], often face such a challenge. Existing methods in rule extraction focus on global explainability, which compromises the regional explainability of minor subgroups of data to achieve better overall performance across the entire data distribution.
@@
Figure 1: An example scenario of applying regional rule extraction for diabetes prediction. The data region of interest is the region that contains samples predicted as positive diabetes cases by the model. The results are from our experiments on a diabetes prediction task [10]. In the extracted rules, Hemoglobin A1c (HbA1c) is a measure of a person’s average blood sugar level over the past 2-3 months.
@@
Figure 2: The workflow of our methods. Algorithm 1 is optional, depending on the dimensionality of the data.
@@
To better interpret data patterns for a specified subgroup, regional rule extraction aims to extract optimal rules from a specified data region. Formally, the outcome of regional rule extraction from a model is in the form of IF-THEN statements: IF $x\in\mathbb{X}$, THEN $y\in\mathbb{Y}$, here $\mathbb{X}$ denotes a subspace of a multivariate variable $x$ confined by a set of rules, e.g. value ranges of several features. $\mathbb{Y}$ denotes a subset of a target variable $y$. For instance, when extracting rules from a classifier, $y$ often refers to the model’s prediction and $\mathbb{Y}$ denotes a class of interest (e.g. positive class in binary classification).
@@
Compared to global rule extraction, regional rule extraction does NOT generate rules for the data region $y\notin\mathbb{Y}$. It is expected to gain two advantages through this reduction: i) obtaining more accurate and generalized rules for the specified data region $y\in\mathbb{Y}$, and ii) providing more precise control on key properties of the extracted rules (e.g. number of rules, number of samples that satisfy these rules).
@@
The general objective of regional rule extraction is to search for the optimal subspace of features that maximally elevates the occurrence of the data of interest. Conventional ways to extract rules from data often utilize various search strategies, such as hill-climbing, beam search, and exhaustive search [11]. Due to the vastness of the search space, these strategies often require predefined discretization for numerical features. However, predefined discretization is neither applicable nor desirable in many real-world applications. And in such a preprocessing, features are often discretised individually because the computational cost may increase exponentially when considering dependencies among a large number of features. Such individual discretization considers only the marginal distributions of features rather than their joint distributions, there is an implicit assumption that these features are independent – a condition not valid in many real-world scenarios.
@@
    we enhance regional rule extraction for black-box machine learning models, offering more accurate and generalized rules for underrepresented data regions than global rule extraction, and also allowing for local rule extraction for a given sample;
@@
    we enable automatic and adaptive rule generation with numerical features by integrating the discretization process into the search strategy for a feature subspace, without requiring predefined discretization or making any assumption about feature distributions;
@@
The workflow of our methods is illustrated in Figure 2. To the best of our knowledge, this is the first attempt to specifically address regional rule extraction with automated rule generation in the literature of XAI.
II Related Work
Rule extraction is the most human-friendly solution in global XAI for knowledge-intensive applications because it can directly provide human-understandable explanations as long as the data features are interpretable. This property is particular desirable for tabular data, which is prevalent in domains such as healthcare, finance, and social science. Nonetheless, most efforts in global XAI have been focused on computer vision tasks [12, 13, 14] and language processing tasks [15, 16, 17], which primarily rely on visualization and semantic concepts to bring interpretability to deep models. To improve regional explainability for various types of data, we aim to develop a general method of rule extraction for identifying regional data patterns.
There are three conceptual categories of the implementation of rule extraction in neural networks [3, 4]: decompositional methods [18, 19, 20, 21] – the rule extraction is performed layer by layer and involves every neuron in a model; pedagogical methods [22, 23, 24] – where only the input and output of a model are involved in the rule extraction; eclectic methods [25, 26] – the rule extraction can be performed through a combination of decompositional and pedagogical approaches, involving the grouping of components within a neural network. In this work, our focus is on pedagogical rule extraction which is model-agnostic.
One main challenge of rule extraction is the discretization of numerical features, with the goal of locating the most informative value intervals for numerical features. Most existing methods require pre-defined discretization to reduce the search space. For instance, association rule classifiers [27, 28] use association rule mining [29, 30] to transform rule extraction into finding frequent item sets – groups of items that frequently appear together in training samples. Consequently, numerical features must be transformed to discretized items before this process. Similarly, KDRuleEx [23] was developed to extract decision tables using categorical features. MUSE [31] is a method for obtaining faithful rule sets in a feature subspace, where the candidate rules of numerical features must be readily provided. Moreover, Falling rule list [32], Scalable Bayesian Rule Lists (SBRL) [33], and Bayesian Rule Sets (BRS) [34] are probabilistic rule extraction methods that use distributions of discrete variables for modeling rule generation.
A few prior work made attempts to automate the discretization process. For example, REFNE [35] utilizes ChiMerge [36] to automatically discretize numerical features for extracting rules from model ensembles. It relies on an augmented data set that is generated by model ensembles to enable ChiMerge as an automated process for rule generation. In addition, it ignores interactions between features because each feature is discretized independently.
Among the existing methods with automated rule generation, Decision Tree (DT) [37, 38] is the most popular one applied in rule extraction for neural networks due to its flexibility and accessibility [18, 19, 20]. DTs are inherently pedagogical, yet they are often utilized in decompositional methods to construct decision paths between intermediate layers. Each decision path to a leaf node in a decision tree naturally represents a rule set associated with a data subgroup. However, DT is not well-suited for regional rule extraction because the criteria of splitting (segmenting) numerical features are based on a global objective that usually compromises performance on minor subgroups. In our experiments, we found that up-weighting the minor class in DT can not improve it’s performance on regional rule extraction. We posit this is because the selection of feature splits may still be affected by other data regions that are not of interest, especially in multi-mode scenarios.
A relevant concept of regional rule extraction is sequential covering, which iteratively learns a rule list for a specific data subgroup until a decision set that covers the entire dataset is composed [11]. In each step of sequential covering, learning the rule list for a specified subgroup is a typical scenario of regional rule extraction.
III Methods
In the rest of this article, we use following notations: $\mathbb{X}$ denotes a rule set with only conjunctive rules, $|\mathbb{X}|$ denotes the length of the rule set, $\mathbb{Y}$ denotes a specified value subset of the target variable $y$, $n$ is the sample index, $N$ is the number of training samples, and $\mathbb{I}(\cdot)$ denotes the indicator function.
Formally, the objective of regional rule extraction is to find a subspace of data features that maximizes the purity of a specified data subgroup within it. The feature subspace is defined by a rule set with two constraints: i) $l_{max}$ – the maximum number of rules included in the rule set, constraining the complexity of the rule set to not be too large; ii) $s_{min}$ – the minimum number of training samples that satisfy the rule set, constraining the support of the rules to not be too small. We quantify the purity of a specified data subgroup as its conditional probability within a subspace, then the objective can be formulated as:
We define one rule by one feature, i.e. one value of a categorical feature or one value interval of a numerical feature. Thus, the length of a rule set is also the number of included features.
It is computationally prohibitive to solve such an optimization problem exactly, especially for numerical features in high-dimensional space. The challenges are mainly due to: i) large search space for selecting feature combinations to compose rules; ii) large search space for selecting value intervals of numerical features. We tackle these challenges heuristically through two procedures: i) a feature selection procedure, and ii) a rule extraction procedure that integrates rule generation and selection processes.
(b) Illustration of generating feature intervals in a multi-mode scenario. We can capture both modes by initializing two intervals with the two peak grids.
Figure 3: Visualization of the threshold selection and feature interval searching in a multi-mode scenario.
III-A Feature selection for rule extraction
Although the feature selection procedure is optional, we recommend conducting it before rule extraction for high-dimensional data. To align with the purpose of rule extraction, we focus on selecting features that frequently appear in a decision path, aligning with the principles of decision-making. This can be achieved by mining feature sets that frequently have high importance across a sample set.
We propose a new method (Algorithm 1) for effectively and automatically selecting features of rules. Meanwhile, we introduce a new feature importance measurement (Equation 3) for differentiable models, leveraging integrated gradients [39] to quantify feature importance as it can better align with the principle of decision rules. For non-differentiable models, such as tree models or ensemble models, one can use any feature importance measures that are applicable, e.g., SHAP [40] or LIME [41].
III-A1 Preliminary
The integrated gradients (also known as integrated Jacobian) [39, 42] is a pairwise measure that can quantify the contribution of changing an input feature on the shift of a target variable given by a trained differentiable model. It can be applied to identify feature sets that most frequently have a joint high impact on shifting a target variable. Such feature sets are crucial for determining the target variable’s value.
III-A2 Feature importance measured by integrated gradients
$\hat{\mathbf{j}}_{i}=\left|\frac{\mathbf{j}_{i}}{y-\tilde{y}}\right|,\quad\text{where}\ \ y=\mathcal{G}(x),\ \ \tilde{y}=\mathcal{G}(\tilde{x})$ (3)
In brief, $\hat{\mathbf{j}}_{i}$ represents the ratio of the shift attributed to the $i$-th feature out of the total shift caused by all features.
As the original integrated gradients are sensitive to the scale of the target variable, we normalize the integrated gradients by the absolute value of the shift of the target variable. This normalization ensures that the importance of each feature is comparable across different sample pairs, which is crucial for identifying frequently important feature sets.
Since this feature importance is measured using the shift between a baseline sample and a test sample, we suggest choosing class centers as baseline samples in classification tasks. Test samples should be a randomly selected subset from the training set to simulate decisions about the class origin of a random sample. We ensure the test samples are balanced across all classes, preventing the oversight of important features for minor subgroups of data . For regression or unsupervised tasks, the selection of baseline samples could depend on the specific application. A few general options are the mean, median, or mode of the training samples clustered by target subgroups.
III-A3 Frequently important feature sets
The proposed feature importance (Equation 3) is measured by a baseline sample and a test sample. When given $B$ baseline samples, $M$ test samples, and $D$ features for each sample, one can produce an importance matrix $\mathbf{J}\in\mathbb{R}^{B\times M\times D}$. For instance, $\mathbf{J}_{b,m,i}$ is the importance of the $i$-th feature on the shift of the target variable from the $b$-th baseline sample to the $m$-th test sample. This importance matrix $\mathbf{J}$ can be readily transformed to a size of $BM\times D$. For feature importance determined by other measures, one may directly construct the importance matrix of size $N\times D$, where $N$ is the number of samples.
Then each row of the importance matrix $\mathbf{J}$ can be transformed into a sequence of feature indices, containing indices of features with an importance score greater than a threshold $\mathbf{j}_{th}$. Through this transformation, we are able to apply the FP-Growth algorithm [43] to discover subsets of features that frequently have a high impact together. FP-Growth is an efficient algorithm for mining frequent item sets (i.e. items frequently appear in the same records). Here we treat a feature index as an item and one sequence as one record.
This equation requires the threshold $\mathbf{j}_{th}$ to be sufficiently large so that only one feature can have an impact score larger than $\mathbf{j}_{th}$ in most rows of the importance matrix $\mathbf{J}$. $\gamma BM$ specifies the minimum coverage of the most frequent feature, where $\gamma\in(c_{min}/BM,1]$, and $c_{min}$ is the minimum occurrence of a frequent set. For instance, when $\gamma=1$, it requires that the most frequent feature has impact scores larger than $\mathbf{j}_{th}$ for all samples, similar to the root node of a decision tree. We set $\gamma=0.99$ in all of our experiments. When $\gamma$ is smaller, $\mathbf{j}_{th}$ will be higher, and more features will be filtered out. Our implementation uses a stepwise scanning process to find the threshold. As illustrated in Figure 3a, we initialize the threshold with a small value and increase it stepwise until only one feature has a frequency no less than $\gamma BM$ (Equation 4). It is clear to see that the number of features that satisfy the $\gamma BM$ frequency requirement decreases as the threshold increases. In this way, an appropriate threshold can be automatically selected to include sufficiently representative features for rule extraction. The pseudo-code of feature selection is described in Algorithm 1.
In FP-Growth algorithm, a frequent item set should have appeared at least $c_{min}$ times among all records. Additionally, it is often required to include fewer than $k_{max}$ items. This requires $c_{min}$ and $k_{max}$ as two hyperparameters for the FP-Growth algorithm. $c_{min}$ determines the least frequent item in a frequent item set, so we suggest a relatively small value for it, such as 10% of the training size. The use of threshold $\mathbf{j}_{th}$ automatically limits the number of items in a frequent item set, so we just set $k_{max}$ to be large enough to encompass the longest item set. When there are multiple frequent item sets discovered, we choose the longest one, and if there are ties, we choose the most frequent one as the candidate feature set for rule extraction.
Algorithm 1 FrequentFeaturesSelection
Return: $F^{*}_{fq}$
III-B Automatic and Model-agnostic Regional Rule Extraction (AMORE)
Figure 4: Demonstration of the automatic rule generation with numerical features. This figure was plotted with synthetic data of two classes in a 2-dimensional feature space. The target subgroup is set as $y=1$. We generate the first rule by identifying the mode interval of the probability ratio on the feature $x_{1}$ (the vertical axis). The second rule for the feature $x_{2}$ (horizontal axis) is generated by conditioning on the first rule. We applied KDE for plotting density curves, the histograms of probability ratios and the subspace in the red rectangle were obtained through our method. More details can be found in Algorithms 4 and 5.
In general, regional rule extraction is to provide a subset of features with a length $\leq l_{max}$ and the value ranges of those features that can maximize the conditional probability in Equation 1. We can explicitly decompose the rule set to each selected feature, i.e. $\{x\in\mathbb{X}\}\triangleq\{x_{f_{1}}\in\mathbb{X}_{f_{1}},\dots,x_{f_{l}}\in\mathbb{X}_{f_{l}}\}$, where $\{f_{1},\dots,f_{l}\}$ are the indices of selected features, $\mathbb{X}_{f_{i}}$ $(\forall 1\leq i\leq l)$ denotes a value range of $f_{i}$-th feature. According to the chain rule of conditional probability, we can then approximate the conditional probability as below:
$\mathcal{X}_{l}^{i-1}$ denotes conditions composed by previous $i-1$ features and $\mathcal{X}_{l}^{0}$ is the domain of $x$. We treat this decomposition as a sequence of operations. In the $i$-th operation, we select the value range of $f_{i}$-th feature that maximizes a probability ratio $\mathbf{r}_{f_{i}}$, where $\mathbf{r}_{f_{i}}$ is defined as the following:
Through this decomposed objective for each feature, we can obtain candidate rules one by one for a given order of features. However, the result is not guaranteed to be globally optimal if we do not search all possible orders of features. As it is usually computationally prohibitive to find the exact global optimal solution, a feasible way is to select the feature $f_{i}$ that has the maximum $\mathbf{r}_{f_{i}}$ among all remaining features at the $i$-th step, which is analogous to the strategy of adding branches in decision trees. This approach can achieve a global optimal solution when the features are independent of each other, as the order of features does not matter in such cases.
We provide detailed pseudo code of this process in Algorithms 2 and 3. The input argument $n_{g}$ is the number of grids for initializing value intervals of numerical features, which will be explained in Algorithm 4. $K$ is the maximum number of candidate rules that can be added at each step. We may obtain multiple rule sets when $K>1$ to cope with multi-mode scenarios in data distributions. Note that there may be fewer than $K$ rules added at each step as we discard rules with a ratio not larger than 1. We set $K=3$ in all of our experiments.
Algorithm 2 ExtractRuleSets
Input: $X$ – features of training samples; ${Y}$ – indicator of the target subgroup $y\in\mathbb{Y}$, which is a binary vector; $F^{*}_{fq}$ – the selected feature set; $s_{min}$ – minimum support of a rule set; $l_{max}$ – maximum length of a rule set; $n_{g}$ – number of value grids for numerical features; $K$ – maximum number of candidate rules added at each step.
# Start adding rules from the root node, see Algorithm 3.
Return: $\mathbb{R}_{sets}$
Algorithm 3 AddRules
Input: $\bm{\nu}_{par}$ – parent rule node; $F^{*}_{fq}$ – the selected feature set; $X$ – features of training samples; ${Y}$ – indicator of the target subgroup $y\in\mathbb{Y}$, which is a binary vector; $\mathcal{X}_{pre}$ – previous rules; $s_{min}$ – minimum support of a rule set; $l_{max}$ – maximum length of a rule set; $n_{g}$ – number of value grids for numerical features; $K$ – maximum number of candidate rules added at each step.
if $\tilde{F}^{*}_{fq}$ is not empty then
    for $\mathbb{X}^{*}_{f_{i}}\in\mathbb{X}^{*}_{K}$ do
         $\bm{\nu}_{new}\leftarrow$ Create a new rule node by $\mathbb{X}^{*}_{f_{i}}$,
         if $len(\tilde{\mathcal{X}}_{pre})<l_{max}$ then
         end if
    end for
end if
III-B1 Generating rules for numerical features
To generate candidate rules for categorical features, each category can be considered as a candidate for the designated rule. However, it is not trivial to generate candidate rules for numerical features. The question here is how to automatically identify intervals of numerical features that have high enough probability ratios and large enough support. We propose a histogram-based approach that initiates a value interval by a specified grid and iteratively expands the interval to its neighbor grids when certain criteria are met.
This process begins with a binning step that divides the value range of a feature into a number of grids $n_{g}$. We offer a few options for the binning strategy in our implementation, including “uniform”,“kmeans”, and “quantile”. The “uniform” strategy ensures all initial grids in each feature have identical widths; “kmeans” assigns values in each grid have the same nearest centre of a 1D k-means cluster; “quantile” results in grids with an equal number of points across each feature. We found that the “uniform” and “kmeans” strategies work better than “quantile” in all of our experiments.
Intuitively, we should start with a peak grid that gives the highest ratio $\mathbf{r}_{f_{i}}$ and expand it when possible. Nonetheless, this might not be able to find the optimal interval if the ratios in its neighbor grids decrease rapidly. An example is demonstrated in Figure 3b, in which the peak grid in the interval $\mathbb{X}_{a}$ is higher than the peak grid in the interval $\mathbb{X}_{b}$, however, the overall ratio of the interval $\mathbb{X}_{a}$ is lower than the overall ratio of the interval $\mathbb{X}_{b}$. It is also possible that there exist multiple intervals with similar ratios and we can not know which one is better before finishing the search for a whole rule set. Considering such multi-mode scenarios, we search intervals that start with every peak grid and choose candidate intervals with the top $K$ ratios. We implemented a simple method based on first-order derivatives to find peaks of the ratio histogram, which can be replaced by more advanced methods. We only select peak grids with a ratio larger than 1. Here $K$ is the same as the maximum number of candidate rules can be added in each step (Algorithm 3).
Algorithm 4 GetCandidateRules
Input: $X$ – features of training samples; $f$ – index of the specific feature; ${Y}$ – indicator of the target subgroup $y\in\mathbb{Y}$, which is a binary vector; $\mathcal{X}_{pre}$ – previous rules; $s_{min}$ – minimum support of a rule set; $n_{g}$ – number of value grids for numerical features; $K$ – maximum number of candidate rules.
$\mathbf{x}^{f}_{g}\leftarrow$ get $n_{g}$ value grids of $f$-th feature
$\mathbf{s}_{g}\leftarrow$ Calculate supports of each grids,
Return: $K$ rules in $\mathbb{X}_{f}$ with top $K$ ratios in $\mathbf{r}_{f}$
  1. i)
    when the support of this interval is smaller than the minimum support $s_{min}$ we expand the interval to its neighbor grids if possible;
  2. ii)
    when the support of this interval is not smaller than $s_{min}$, we only expand the interval to a neighbor grid when this neighbor grid has a higher ratio. This will guarantee that an interval with enough support only expands when its ratio gets increased (Proposition 1).
The pseudo-code of this rule generation process for numerical features are described in Algorithms 4 and 5. Figure 4 visualizes this process with synthetic data in a 2-dimensional feature space.
Proposition 1.
Proof.
Algorithm 5 GenFeatureInterval
Input: $s_{min}$ – minimum support of a rule set, $g_{p}$ – index of a peak grid to initiate an interval; $\mathbf{x}_{g}$ – value grids of a feature; $\mathbf{r}_{g}$ – probability ratios of value grids; $\mathbf{s}_{g}$ – supports of value grids; ${Y}$ – indicator of the target subgroup $y\in\mathbb{Y}$, which is a binary vector; $\mathcal{X}_{pre}$ – previous rules.
$\mathbb{X}_{f,p},s,r\leftarrow$ Initialize an interval by the peak grid $g_{p}$ and its support $\mathbf{s}_{g}[g_{p}]$, ratio $\mathbf{r}_{g}[g_{p}]$,
while $s<s_{min}$ or $r>1$ do
    if $s<s_{min}$ then
         $\mathbb{X}_{f,p},s,r\leftarrow$ Expand the interval to a neighbor grid that has a larger ratio than the other neighbor grid,
    else if $r>1$ then
         $\mathbb{X}_{f,p},s,r\leftarrow$ Expand the interval to a neighbor grid that has a higher ratio than the other neighbor grid and also higher than $r$. If no such grid exists, break.
    end if
end while
if $s<s_{min}$ or $r\leq 1$ then
    $\mathbb{X}_{f,p}=\emptyset$
end if
Return: $\mathbb{X}_{f,p},r,s$
Additionally, we provide local rule extraction for a given sample (as demonstrated in Table II in Results section), which forces the value interval of a feature to include the feature value of a given sample. If the feature intervals obtained by Algorithms 4 and 5 do not meet this requirement, we apply a local searching process. This can be directly achieved by starting with the grid that includes the given sample. The local rule extraction may not be able to obtain any valid rules (i.e. the ratios of any value intervals are not greater than 1) if the given sample is an outlier of the target subgroup.
III-C Evaluation criteria
We apply three criteria for evaluating the quality of an extracted rule set $\mathbb{X}$ from a model. As we conduct experiments on classification tasks, we denote the true label of a sample as $y^{*}$, the model’s prediction as $\hat{y}$, the class of interest as the target class $c$, the target subgroup for rule extraction is given by the model’s prediction as $\hat{y}=c$.
  1. 1.
    $\text{Support}=\sum_{n=1}^{N}\mathbb{I}(x^{(n)}\in\mathbb{X})$
  2. 2.
    Confidence (also called rule accuracy) [28] – the proportion of samples that belong to the target subgroup while satisfying the rule set. It measures the purity of the target subgroup given the rule set:
  3. 3.
    Fitness – a new metric we propose for regional rule extraction, being the ratio difference between samples that belong to the target subgroup and those that do not, under the rule set, divided by the total number of samples in the target subgroup. The higher the score, the better, and in the best case $\text{Fitness}=1$:
Support and confidence are widely applied in the literature of rule extraction. Typically, support and confidence are negatively correlated because including more samples likely introduces additional noise into the subspace. A better rule set should exhibit reduced correlation by increasing confidence while maintaining a similar support. We define fitness for evaluating regional rule sets, which takes into account the trade-off between support and confidence. This metric measures how well the subspace defined by a rule set fits into the target data region. A rule set with larger support and lower confidence, or one with smaller support and higher confidence, could both result in a lower fitness. Fitness can be negative when the subspace is mostly out of the target subgroup.
IV Results